Having a little with this equation. The length of a rectangle is 5in longer than its width. If the perimeter of the rectangle is 42in finds its area.
Answer by ikleyn(52781)

.
2L + 2W = 42


L + W = 42/2 = 21


W + (W+5) = 21


2W = 21 - 5 = 16  ====>  W = 16/2 = 8.


Width = 8 inches.


Length = 8 + 5 = 13 inches.


Area = L*W = 8*13 = 104 square inches.